OpenStudy (anonymous):

I'm looking for the first derivative of f(x)=csc^2(sqrtx) basically how do i get rid of the csc^2

OpenStudy (anonymous):

can i do csc x csc(sqrtx) ?

OpenStudy (anonymous):

let put that better csc(csc(sqrt(x)))

OpenStudy (anonymous):

You have to use the chain rule. The derivative of csc(x)=-csc(x) cot(x) \[\csc(x^{\frac{1}{2}})(-\csc(x^{\frac{1}{2}}) \cot(x^{\frac{1}{2}}))(\frac{1}{2}x^{-\frac{1}{2}})\]

OpenStudy (anonymous):

what is that though the multiplication rule?!?!

OpenStudy (anonymous):

Through the chain rule. First you deal with f(____)=____^2

OpenStudy (anonymous):

csc^2

OpenStudy (anonymous):

Oh, shoot, I forgot the 2.

OpenStudy (anonymous):

It should be 2csc___

OpenStudy (anonymous):

ok i see thats just the derivative of csc^2 right?

OpenStudy (anonymous):

Right. And you leave the inside the same. Then you take the derivative of csc next.

OpenStudy (anonymous):

but why do you take the derivative twice there? the chain rule kind of confuses me...

OpenStudy (anonymous):

Think of it this way. You have a composition of functions: \[f(x)=x^2\]\[g(x)=\csc(x)\]\[h(x)=x^{\frac{1}{2}}\] When you combine them, you get \[f(g(h(x)))=(\csc(x^{\frac{1}{2}}))^2\]

OpenStudy (anonymous):

That's what your original function was. All you have to do is take the derivative of each function individually. The first thing we did was take care of f(x). We have to take the derivative of csc next because that's g(x)

OpenStudy (anonymous):

so, f(s)=2x g(x)=-csc(x)cot(x) h(x)=1/2x^-1/2

OpenStudy (anonymous):

so -2xcsc(1/2x^-1/2)cot(1/2x^-1/2)

OpenStudy (anonymous):

and thats my final answer?!?!

OpenStudy (anonymous):

Close. The only catch is that you're not taking f(x). You're taking\[f(\csc(x^{\frac{1}{2}})\] That means that you don't have 2(x). You have \[2(\csc(x^{\frac{1}{2}}))\] Later on, you're not taking g(x). You're taking \[g(x^{\frac{1}{2}})\] That means you don't have csc(x)cot(x). You have \[\csc(x^{\frac{1}{2}})\cot(x^{\frac{1}{2}})\]

OpenStudy (anonymous):

Forgot a - in the csc part

OpenStudy (anonymous):

ok i can work with that

OpenStudy (anonymous):

so just add negative sign up there?

OpenStudy (anonymous):

what what about the 1/2x^-1/2?

OpenStudy (anonymous):

Final answer should look like \[2\csc(x^{\frac{1}{2}})(-\csc(x^{\frac{1}{2}})\cot(x^{\frac{1}{2}}))(\frac{1}{2}x^{-\frac{1}{2}})\] That simplifies to \[-\frac{1}{x^2} \csc^2 (x^{\frac{1}{2}})\cot(x^{\frac{1}{2}})\]

OpenStudy (anonymous):

The part with the 1/2x^(-1/2) was good because you were actually taking h(x).

OpenStudy (anonymous):

but why isn't that in our final answer?

OpenStudy (anonymous):

Which one? The one I wrote up above was an unsimplified version.

OpenStudy (anonymous):

the 1/2x^-1/2, and my online homework said that the simplified answer up there was incorrect

OpenStudy (anonymous):

I'm pretty sure it's correct. Maybe it wants the less simplified version? What did you enter?

OpenStudy (anonymous):

the simplified let try the less simplified one

OpenStudy (anonymous):

apparently that was it man that did it.

OpenStudy (anonymous):

Oh good. It's unfortunate that it wouldn't accept a more simplified version.